OpenStudy (anonymous):

When an opinion poll calls residential telephone numbers at random, only 20% of the calls reach a live person. You watch the random digit dialing machine make 15 calls. (a) What is the probability that exactly 5 calls reach a person?

OpenStudy (anonymous):

Use binomial distribution\[P(k)=\left(\begin{matrix}n \\ k\end{matrix}\right)p^k(1-p)^{n-k}\] where n=15, k=5 and p=0.2

OpenStudy (anonymous):

.1032 use the binomial distribution table